OpenStudy (anonymous):

Isadora bought 4 trees to plant in her backyard and spent no less than eighty-four dollars altogether. How many dollars could each tree have cost? Let t represent how many dollars each tree cost.

OpenStudy (anonymous):

Well.... \[ t = \frac{\text{total}}{\text{number_of_trees}} = \frac{$84}{4} = $21 \] each tree 'could' cost $21

OpenStudy (anonymous):

is it greater than or less than

OpenStudy (anonymous):

Well. since it's minimum eighty four then \[ t \ge \frac{$84}{4} \\ t \ge $21 \] Sorry, I was too tired yesterday to pay attention. Missed the "no less than"

OpenStudy (anonymous):

thanx